# Towards Goal-Oriented Semantic Communications: New Metrics, Framework, and Open Challenges

### v2

@@ V Case Study: Goal-oriented Sparse Semantic Sampling
This section considers the real-time wireless fire monitor and rescue system in Fig. 3 as a case study. The goal in Fig. 3 is to minimizing the long-term average economic losses and resource consumption associated with fire and rescue induced by fire and rescue. As an initial exploration, our focus in this article is on the design of sparse semantics-aware sampler. The sampler should decide which semantics to sample and transmit, such that the rescue is timely and effective to minimize the long-term real-world costs induced by fire and rescue. This E2E status mismatch cost could be recorded by GoT, and thus to achieve the goal is equivalent to minimizing the long-term average E2E status mismatch costs.
@@
[…]
@@
$\bullet$ Optimal GoT. Sampling is designed to minimize long-term average E2E mismatch costs recorded by GoT, which is resulted by a combination of personnel casualties, equipment and building damages, as well as manpower or resources consumption in the real world. The sampling policy is solved resorting to the Markov Decision Process (MDP).
Fig. 4 showcases the simulated results that compare the performance of the discussed sampling policies. The performance is evaluated by two indicators: long-term average cost, reflecting the cumulative real-world implications of fire and rescue, and average sample rate, indicating communication overhead. The primary goal is to reduce the long-term average cost, and as Fig. 4 reveals, the GoT-optimal policy excels in this, marking its inherent alignment with our goal. This policy also demonstrates an efficient balance between reduced communication overhead (sample rate) and effectiveness, suggesting a semantics-aware, goal-driven sampling method. For example, the GoT-optimal policy can adaptively identify the semantics “Big Fire” as highly significant, thereby prioritizing its transmission to the receiver. In summary, the GoT-optimal policy excels by focusing on the transmission of data that is most significant to achieve the goal. This method enhances the recognition of semantic significance.
@@ VI Conclusions and Future Challenges
The significance of information plays a critical role in interpreting ‘semantics’. Following this avenue, this work moves further by proposing a new goal-oriented metric and envisioning a holistic goal-oriented architecture. The proposed GoT provides a unified and extensible solution to characterize the goal of communications, which provides a solution to the effectiveness problem. The proposed goal-oriented network architecture demonstrates its great superiority in substantially alleviating the communication burden of the next-generation Internet of Everything (IoE) networks. A preliminary instantiated case study is demonstrated to address the challenge of sparse semantic sampling presented in [7] and [8].
@@
Goal-oriented physical (PHY) layer techniques. To achieve a specific goal under constrained resources, the bit-by-bit reconstruction is so energy-intensive and low-efficiency that it could not perceive the priority of information in terms of the ultimate effectiveness to adaptively allocate the limited resources. In this regard, new goal-oriented paradigms at the PHY could be explored to further improve communication efficiency. In particular, the goal-oriented channel coding and decoding algorithms, the goal-oriented retransmission and feedback mechanism, the multi-user power allocation mechanism, the goal-oriented modulation and signal shaping are some promising candidates. By this means, the grand vision beyond the traditional paradigm is that the future communication is not designed to only reduce the error probability in an average manner but to circumvent the severe errors, where the severity and the cost of the error is recorded by the GoT.
Goal-oriented Perception, Communication, Computation, and Control Co-design . The life cycle of information is closely affiliated with the processes of perception, communication, computation, caching, and control. Therefore, a true leap forward can be achieved by merging these modules together for the co-design optimizations. Particularly, the multi-modal sensors consecutively perceive high-resolution raw data, the computation leads to precise semantic extractions and effective intelligent decisions, the communication serves as a priority-aware information pipeline for information disseminating, and the actuators executes the control commands precisely, so as to transform the information into practical usage. To achieve this co-design, high-order GoTs might be designed to integrate the entire perception-action closed-loop process across various function modules.

### v3

@@ V Case Study: Goal-oriented Sparse Semantic Sampling
This section considers the real-time wireless fire monitor and rescue system in Fig. 3 as a case study. The goal in Fig. 3 is to minimizing the long-term average economic losses and resource consumption associated with fire and rescue induced by fire and rescue. As an initial exploration, our focus in this article is on the design of sparse semantics-aware sampler. The sampler should decide which semantics to sample and transmit, such that the rescue is timely and effective to minimize the long-term real-world costs induced by fire and rescue. This E2E status mismatch cost could be recorded by GoT, and thus to achieve the goal is equivalent to minimizing the long-term average GoT.
@@
As a case study, we evaluate a basic system where the context status $\Phi(t)$ is constant. The semantic status captures three fire severity levels. In this way, the $3\times 3$ GoT is visualized at the right-hand side of Fig. 3, signifying that the semantics “NO FIRE” estimated as “NO FIRE” will consume zero real-world cost, while “SMALL FIRE” estimated as “NO FIRE” will incur the 20 real-world cost per time slot, etc. We assume perfect semantic quantization. The semantic sampler decides when and what to sample and transmit a status through an $\epsilon$-erasure channel to the fire department. Firefighters (actuators) are dispatched based on the latest fire levels. The time is slotted and the source is modeled as a controlled Markov Source. The transition probabilities of the fire levels depend on the firefighters assigned. In such a system, we compare the following six sampling policies in terms of goal-oriented effectiveness and sampling rate:
@@
$\bullet$ Optimal GoT. Sampling is designed to minimize long-term average GoT, which is equivalent to minimizing the long-term average cost resulted by personnel casualties, equipment and building damages, as well as manpower or resources consumption in the real world. The sampling policy is solved resorting to the Markov Decision Process (MDP).
Fig. 4 showcases the simulated results that compare the performance of the discussed sampling policies. The performance is evaluated by two indicators: long-term average cost, reflecting the cumulative real-world implications of fire and rescue, and average sample rate, indicating communication overhead. The primary goal is to reduce the long-term average cost, and as the Fig. 4 reveals, the GoT-optimal policy excels in this, marking its inherent alignment with our goal. This policy also demonstrates an efficient balance between reduced communication overhead (sample rate) and effectiveness, suggesting a semantics-aware, goal-driven sampling method. For instance, the semantic status “BIG FIRE” is always more important due to the fact that it will incur more severe cost. These sampling policies are all event-triggered; therefore, the average cost and sampling rate are derived concurrently. From Fig. 4, the proposed GoT-optimal policy stands out by emphasizing the most goal-relevant, semantics-rich data while ensuring minimum costs.
In a general case, the most important thing is to tailor the GoT to align the specific scenarios and goals. This process motivates accessing the real-world costs of context-specific status mismatch categories. With the tailored GoT in hand, the goal-relevant sparse sampling could be developed similarly.
@@ VI Conclusions and Future Challenges
The relevance of information plays a critical role in interpreting ‘semantics’. Following this avenue, this work moves further by proposing a new goal-oriented metric and envisioning a holistic goal-oriented architecture. The proposed GoT provides a unified and extensible solution to characterize the ultimate goal of communications, which provides a solution to the effectiveness problem. The proposed goal-oriented network architecture demonstrates its great superiority in substantially alleviating the communication burden of the next-generation Internet of Everything (IoE) networks. A preliminary instantiated case study is demonstrated to address the challenge of sparse semantic sampling presented in [7] and [8].
@@
Goal-oriented physical (PHY) layer techniques. To achieve a specific goal under constrained resources, the bit-by-bit reconstruction is so energy-intensive and low-efficiency that it could not perceive the priority of information in terms of the ultimate effectiveness to adaptively allocate the limited resources. In this regard, new goal-oriented paradigms at the PHY could be explored to further improve communication efficiency. In particular, the goal-oriented channel coding and decoding algorithms, the goal-oriented retransmission and feedback mechanism, the multi-user power allocation mechanism, the goal-oriented modulation and signal shaping are some promising candidates. By this means, the grand vision beyond the traditional paradigm is that the future communication is not designed to only reduce the error probability in an average manner but to circumvent the severe errors.
Goal-oriented Perception, Communication, Computation, and Control Co-design . The life cycle of information is closely affiliated with the processes of perception, communication, computation, caching, and control. Therefore, a true leap forward can be achieved by merging these modules together for the co-design optimizations. Particularly, the multi-modal sensors consecutively perceive high-resolution raw data, the computation leads to precise semantic extractions and effective intelligent decisions, the communication serves as a priority-aware information pipeline for information disseminating, and the actuators executes the control commands precisely, so as to transform the information into practical usage.